# Plasmon-assisted photoelectron emission in a model cluster using time-dependent density functional theory and the time–dependent surface–flux method

### v1

@@ Abstract
We investigate plasmon-assisted photoelectron emission using a one-dimensional time-dependent density-functional theory (TDDFT) model. Photoelectron spectra are computed with the time-dependent surface-flux (t-SURFF) method. In addition to the expected above-threshold ionization (ATI) comb, we observe peaks that arise from long-lived plasmon oscillations and the associated electron emission occurring after the laser pulse. We further analyze the positions of these peaks and their scaling behavior with the laser intensity.
@@ I Introduction
Plasmonic nanostructures concentrate optical energy into deep-subwavelength volumes, enabling strong-field photoemission, ultrafast electron sources, and nanoscale control of light–matter interactions Li et al. (2013); Putnam et al. (2017); Stockman (2011); Dombi et al. (2020); Kim et al. (2023). Ultrafast photoemission electron microscopy (PEEM) has directly imaged plasmon dynamics in space and time, providing unique insight into how localized fields drive electron emission at the nanoscale Dabrowski et al. (2020); Kosar and Dani (2024); Ghosh et al. (2024). On the theoretical side, real-time time-dependent density functional theory (TDDFT) Ullrich (2011) has emerged as a powerful framework for following collective electronic motion, charge transfer, and field-driven dynamics in real time and real space, complementing frequency-domain linear-response TDDFT and classical electrodynamics approaches Uchida and Watanabe (2017); Ma et al. (2015); Herring and Montemore (2023); Wang et al. (2024).
@@
In TDDFT, the interacting electron system and linear TDSE are replaced by a formally exact non-interacting Kohn–Sham (KS) system described by nonlinear time-dependent Kohn–Sham (TDKS) equations. Within this framework, a plasmon corresponds to a collective oscillation of KS particles that all experience the same time-dependent KS potential. The exact KS potential, if known, would reproduce the same dipole expectation value as the full many-body solution. In practice, however, commonly used approximate functionals, such as the adiabatic local-density approximation (ALDA), tend to produce persistent oscillations after the laser pulse, as we will demonstrate in this work. Furthermore, the exact observable for photoelectron emission in TDDFT is unknown Véniard et al. (2003); Brics and Bauer (2013), and it is customary to interpret the KS particles themselves as electrons when computing photoelectron spectra.
Photoelectron spectra (PES) in strong-field physics are commonly obtained using (i) projections onto field-free continuum states, (ii) the window-operator method Schafer and Kulander (1990); Bauer (2017), or (iii) the time-dependent surface-flux method (t-SURFF) Tao and Scrinzi (2012); Bauer (2017); Tulsky and Bauer (2020a). Methods (i) and (ii) are difficult to apply in TDDFT because they require a stationary potential after the laser pulse, whereas the KS potential continues to evolve. Method (iii), t-SURFF, avoids this assumption but requires a sufficiently long post-pulse propagation so that electrons of the lowest kinetic energy of interest have reached the flux surface. For the TDSE, whose Hamiltonian is stationary after the pulse, this post-propagation can be performed in a single numerical step. In TDDFT, however, the KS Hamiltonian remains time-dependent after the pulse, precluding such shortcuts.
In this work, we show that the t-SURFF photoelectron spectra obtained within TDDFT exhibit sharp peaks whose positions and widths depend strongly on the length of the post-pulse propagation. These peaks originate from plasmon-assisted electron emission occurring after the pulse: the longer the post-pulse propagation, the narrower the peaks become. Although this sensitivity is clearly an artifact of approximate exchange–correlation functionals, plasmon-assisted electron emission is a genuine physical effect. This enables us to analyze, for example, the scaling behavior of plasmon-assisted electron emission, while bearing in mind that its magnitude is likely overestimated by TDDFT with simple exchange–correlation potentials.
@@ II Model
We model a one–dimensional (1D) cluster of $N_{e}=40$ electrons confined by $N_{\text{ion}}=N_{e}$ equidistantly spaced ions,
@@
with $X_{i}=(-(N_{\text{ion}}+1)/2+i)a$ and $a=1.125$. We use TDDFT to simulate this system in a laser field with vector potential $A(t)$ in dipole approximation. The TDKS equation for the $N_{\text{KS}}=N_{e}/2$ spin-degenerate KS orbitals reads
@@
with
$V_{\text{KS}}[n](x,t)=V_{\text{ion}}(x)+V_{\text{H}}[n](x,t)+V_{\text{xc}}[n](x,t)$ (3)
where
$V_{\text{H}}[n](x,t)=\int\text{d}x^{\prime}\,\frac{n(x^{\prime},t)}{\sqrt{(x-x^{\prime})^{2}+1}}$ (4)
is the Hartree potential and
$V_{\text{xc}}[n](x,t)\simeq V_{\text{x}}[n](x,t)=-\left(\frac{3}{\pi}n(x,t)\right)^{1/3}$ (5)
is the exchange-correlation potential in exchange-only adiabatic local density approximation. Note that we do not aim at modelling a 1D electron system (such as in a nano wire, for instance) but rather a 3D electron system along the laser polarization direction. This is the reason why we use the expression (5) for a 3D electron gas in our 1D model. The electronic density is
$n(x,t)=2\sum_{i=1}^{N_{\text{KS}}}|\varphi_{i}(x,t)|^{2}.$ (6)
Figure 1 shows the ground state KS potential and the 202020 ground state KS probability densities $|\varphi_{i}(x)|^{2}$, shifted to their respective KS energies $\varepsilon_{i}$.
Figure 1: Kohn–Sham potential $V_{\rm KS}(x)$ (solid black) of the 1D cluster and the 20 occupied KS orbitals (coloured) plotted as $|\varphi_{i}(x)|^{2}+\varepsilon_{i}$.
@@
To identify collective eigenmodes we apply a weak $\delta$–kick in the electric field (i.e., a $\Theta$-step in the vector potential) to the ground state configuration, and Fourier transform the total dipole
The pre-factor $\Omega^{2}$ arises because we are actually interested in the FFT of $\ddot{D}(t)$, the dipole acceleration. In order to identify whether a linear-response peak is due to single-particle transitions or collective in nature it is useful to look at the response of individual orbitals.
We employ the t-SURFF method Tao and Scrinzi (2012); Bauer (2017) to compute photoelectron spectra. In 1D and velocity gauge, the momentum–space amplitudes collected at the left (right) surfaces $x_{L}$ ($x_{R}$) read (for each occupied orbital $\varphi_{i}$)
@@ II.1 Numerical details
We use $N=2000$ grid points with a spacing $\Delta x=0.5$, a time step $\Delta t=0.25$, and complex absorbing boundaries at the edges. The convolution theorem is used to calculate $V_{\text{H}}[n](x,t)$. The ground state configuration according
is calculated using imaginary‑time Crank–Nicolson propagation. Real‑time propagation of (2) also uses the Crank–Nicolson scheme but with a predictor-corrector step and absorbing boundaries Bauer (2017). The t-SURFF surfaces are placed at $x_{L}=0.25\,N\Delta x$ and $x_{R}=0.75\,N\Delta x$. Convergence was verified with respect to box size, time step, position of the t-SURFF surfaces, and the $k$–grid used for the spectra.
@@ III Identification of collective modes
We calculate linear response spectra in order to identify where the collective modes in our model system are. Figure 2 shows results for the “frozen” and the fully dynamic system. Here, “frozen” means that we set $V_{\text{KS}}[n](x,t)$ in (2) to $V_{\text{KS}}[n](x,0)$. As a consequence, the “frozen” linear response shows peaks at positions that correspond to energy differences $\varepsilon_{j}-\varepsilon_{i}$ where $\varepsilon_{i}$ is the respective KS energy of orbital $i$ and $\varepsilon_{j}$ is the KS energy of an excited state that is reachable from state $i$ according to dipole selection rules. The expected positions for such peaks are indicated by vertical dashed lines. The strongest peak is at the energy difference between the highest occupied KS orbital ($\varepsilon_{20}=-0.1709$) and the first unoccupied ($\varepsilon_{21}=-0.0984$), indicated by the first gray dashed line. There is no peak at $\varepsilon_{22}-\varepsilon_{20}$ (second gray dashed line) because of the dipole selection rule, while there is one at $\varepsilon_{23}-\varepsilon_{20}$ (third gray dashed line). The same repeats for the second-highest occupied KS orbital at $\varepsilon_{21}-\varepsilon_{19}$ (first green dashed line, no peak), $\varepsilon_{22}-\varepsilon_{19}$ (second green dashed line, peak present) etc. The yellow dashed lines indicate $\varepsilon_{21}-\varepsilon_{18}$, $\varepsilon_{22}-\varepsilon_{18}$, $\varepsilon_{23}-\varepsilon_{18}$. One might also expect a peak at $\varepsilon_{20}-\varepsilon_{19}$ but it interferes away with a transition from $\varepsilon_{20}$ “down to” $\varepsilon_{19}$ (this is how the Pauli principle is realized in a system of non-interacting identical particles).
Figure 2: Linear–response spectra. Black: full, dynamic response according (2). Blue: response for frozen KS potential $V_{\text{KS}}[n](x,0)$. Vertical dashed lines indicate KS level differences. Two strong peaks at $\omega_{A}=0.106$ and $\omega_{B}=0.156$ appear in the fully dynamic (“unfrozen”) response.
In the full, dynamical response (labeled “unfrozen”) two strong peaks at
@@
are observed. In order to prove that these are collective responses we show orbital-resolved linear response spectra in Fig. 3. In such spectra, one sees all the transitions to other states, which, however, are different for the different KS orbitals. Only in the case of a harmonic-oscillator-like KS potential with equidistant levels would there be only one common response at the harmonic oscillator frequency for all orbitals. However, we do see response peaks in Fig. 3 that are at the same $\Omega$ for all KS orbitals despite the fact that $V_{\text{KS}}[n](x)$ is not a harmonic-oscillator potential. By definition, this is a collective response, as all KS orbitals oscillate in unison with this frequency. The first two of such collective responses are at $\Omega=0.106$ and 0.1560.1560.156.
Figure 3: Orbital–resolved spectra. A strong response of all KS orbitals at a given $\Omega$ indicates a collective mode. Instead, orbital-dependent responses are due to single-particle transitions.
The current density in space and time for those two collective modes are shown in Figure 4. The oscillations are plotted for times after the laser pulses that excited these modes selectively with laser frequencies $\omega_{L}=\omega_{A}/3$ and $\omega_{L}=\omega_{B}/3$, respectively. As expected, the lower-energetic mode $\omega_{A}$ amounts to the oscillation of the electron density as a whole. The higher-energetic mode displays two nodes close to the edges of the 1D cluster. The current density inside the cluster is thus opposite to the current densities at the “surface”. We will use in the following “collective mode” and “plasmon” synonymously.
Figure 4: Oscillating current densities for the collective modes $\omega_{A}=0.106$ (upper panel) and $\omega_{B}=0.156$ (lower panel). The modes were excited with laser pulses of frequencies $\omega_{L}=\omega_{A}/3$ and $\omega_{L}=\omega_{B}/3$, respectively. The oscillating current densities are shown for times after the laser pulses.
@@ IV Photoelectron spectra
Figure 5 shows the photoelectron spectrum (12) for a laser pulse with parameters $a_{0}=0.004$, $\omega_{L}=0.052$, and $N_{\text{cyc}}=20$. Also shown are the individual spectra for the highest occupied KS orbitals $i=20,19,18$. Lower-lying KS orbitals do not play a role for the total electron yield. Besides the expected ATI peaks, separated by $\omega_{L}$, there are very sharp spiky features. The more detailed analysis presented below shows that these features arise from plasmon excitations that decay too slowly in the TDDFT treatment. Extended post-pulse propagation within the t-SURFF framework is required to collect low-energy electrons at the distant t-SURFF surfaces. During this propagation, t-SURFF also accumulates probability amplitude associated with plasmon-assisted electron emission occurring after the laser pulse. The longer the post-pulse duration, the more pronounced and sharper these peaks become.
@@
Figure 5: Total photoelectron spectrum (red) and the individual contributions from the three highest initially populated KS orbitals for a pulse that dominantly excites $\omega_{B}=0.156$ with three laser photons. Narrow spikes are equally spaced by $\omega_{B}$. Laser parameters are $a_{0}=0.004$, $\omega_{L}=0.052$, $N_{\text{cyc}}=20$.
@@ IV.1 Time-of-flight analysis
Figure 6: Windowed t-SURFF spectrum. Laser parameters are the same as in Fig. 5, i.e., $a_{0}=0.004$, $\omega_{L}=0.052$, $N_{\text{cyc}}=20$.
Figure 6 shows the windowed t-SURFF spectrum for thee same laser parameters as in Fig. 5. The width of the Gaussian time-window was one laser cycle so that the ATI peaks due to direct, laser-induced ionization are clearly visible. All ATI-emitted electrons arrive before $\omega_{L}t/2\pi\simeq 30$ at the t-SURFF surfaces, which is consistent with them being emitted during the laser pulse plus their time-of-flight. After the ATI peaks, the plasmon-assisted emission builds up and is clearly visible as horizontal stripes at positions $E^{(A,B)}_{in}=\varepsilon_{i}+n\,\omega_{A,B}$. The stripes are broader in energy than the narrow spikes in the total spectrum in Fig. 5 because of the narrow time-window. This also shows that the plasmon-assisted electron emission peaks would be broader (and probably closer to reality) if we stop the simulation soon after the laser pulse. However, at that time the low-energetic electrons have not yet arrived at the t-SURFF surfaces and thus are missing in the spectra. We also experimented with “freezing” the KS potential after the laser pulse, which, however, yields other artefacts.

### v2

# Plasmon-assisted photoelectron emission in a model cluster using time-dependent density functional theory and the time-dependent surface-flux method

@@ Abstract
We investigate plasmon-assisted photoelectron emission using a one-dimensional time-dependent density-functional theory (TDDFT) model. The plasmons are excited nonlinearly by three laser photons. Photoelectron spectra are computed with the time-dependent surface-flux (t-SURFF) method. In addition to the expected above-threshold ionization (ATI) comb, we observe peaks that arise from long-lived plasmon oscillations and the associated electron emission occurring after the laser pulse. We further analyze the positions of these peaks and their scaling behavior with the laser intensity.
@@ I Introduction
Plasmonic nanostructures concentrate optical energy into deep-subwavelength volumes, enabling strong-field photoemission, ultrafast electron sources, and nanoscale control of light-matter interactions Li et al. (2013); Putnam et al. (2017); Stockman (2011); Dombi et al. (2020); Kim et al. (2023). Ultrafast photoemission electron microscopy (PEEM) has directly imaged plasmon dynamics in space and time, providing unique insight into how localized fields drive electron emission at the nanoscale Lemke et al. (2013); Dabrowski et al. (2020); Kosar and Dani (2024); Ghosh et al. (2024). On the theoretical side, real-time time-dependent density functional theory (TDDFT) Ullrich (2011) has emerged as a powerful framework for following collective electronic motion, charge transfer, and field-driven dynamics in real time and real space, complementing frequency-domain linear-response TDDFT and classical electrodynamics approaches Uchida and Watanabe (2017); Ma et al. (2015); Herring and Montemore (2023); Wang et al. (2024).
@@
In TDDFT, the interacting electron system and linear TDSE are replaced by a formally exact non-interacting Kohn-Sham (KS) system described by nonlinear time-dependent Kohn-Sham (TDKS) equations. Within this framework, a plasmon corresponds to a collective oscillation of KS particles that all experience the same time-dependent KS potential. The exact KS potential, if known, would reproduce the same dipole expectation value as the full many-body solution. In practice, however, commonly used approximate functionals, such as the adiabatic local-density approximation (ALDA), tend to produce persistent oscillations after the laser pulse, as we will demonstrate in this work. Furthermore, the exact observable for photoelectron emission in TDDFT is unknown Véniard et al. (2003); Brics and Bauer (2013), and it is customary to interpret the KS particles themselves as electrons when computing photoelectron spectra.
Photoelectron spectra (PES) in strong-field physics are commonly obtained using (i) projections onto field-free continuum states, (ii) the window-operator method Schafer and Kulander (1990); Bauer (2017), or (iii) the time-dependent surface-flux method (t-SURFF) Tao and Scrinzi (2012); Bauer (2017); Tulsky and Bauer (2020a). Methods (i) and (ii) are difficult to apply to TDDFT because they require a stationary potential after the laser pulse, whereas the KS potential continues to evolve. Method (iii), t-SURFF, avoids this assumption but requires a sufficiently long post-pulse propagation so that electrons of the lowest kinetic energy of interest have reached the flux surface. For the TDSE, whose Hamiltonian is stationary after the pulse, this post-propagation can be performed in a single numerical step. In TDDFT, however, the KS Hamiltonian remains time-dependent after the pulse, precluding such a shortcut.
In this work, we show that the t-SURFF photoelectron spectra obtained within TDDFT exhibit sharp peaks whose positions and widths depend on the length of the post-pulse propagation. These peaks originate from plasmon-assisted electron emission occurring after the pulse: the longer the post-pulse propagation, the narrower the peaks become. Although this sensitivity is clearly an artifact of approximate exchange-correlation functionals, plasmon-assisted electron emission is a genuine physical effect. This enables us to analyze, for example, the scaling behavior of plasmon-assisted electron emission, while bearing in mind that its magnitude is likely overestimated by TDDFT with simple exchange-correlation potentials.
Atomic units are used unless noted otherwise.
@@ II Model
We model a one-dimensional (1D) cluster of $N_{e}=40$ electrons confined by $N_{\text{ion}}=N_{e}$ equidistantly spaced ions,
@@
with $X_{i}=(-(N_{\text{ion}}+1)/2+i)a$ and $a=1.125$. The value for $a$ was chosen such that the KS energy for the valence orbital is close to the work function of silver clusters.
Density functional theory (DFT) is used to calculate the ground state configuration for $N_{\text{KS}}=N_{e}/2$ spin-degenerate KS orbitals according to the KS equation
with
$V_{\text{KS}}[n](x)=V_{\text{ion}}(x)+V_{\text{H}}[n](x)+V_{\text{xc}}[n](x)$ (3)
where
$V_{\text{H}}[n](x)=\int\text{d}x^{\prime}\,\frac{n(x^{\prime})}{\sqrt{(x-x^{\prime})^{2}+1}}$ (4)
is the Hartree potential and
$V_{\text{xc}}[n](x)\simeq V_{\text{x}}[n](x)=-\left(\frac{3}{\pi}n(x)\right)^{1/3}$ (5)
is the exchange-correlation potential in exchange-only adiabatic local density approximation. Note that we do not aim at modelling a 1D electron system (such as in a nano wire, for instance) but rather a 3D electron system along the laser polarization direction. This is the reason why we use the expression (5) for a 3D electron gas in our 1D model.
We use TDDFT to simulate this system in a laser field with vector potential $A(t)$ in dipole approximation. The TDKS equation reads
@@
where $V_{\text{KS}}[n](x,t)$ corresponds to (3) with the time-dependent density
$n(x,t)=2\sum_{i=1}^{N_{\text{KS}}}|\varphi_{i}(x,t)|^{2}$ (7)
inserted. Figure 1 shows the ground state KS potential and the $20$ ground state KS probability densities $|\varphi_{i}(x)|^{2}$, shifted to their respective KS energies $\varepsilon_{i}$.
Figure 1: Kohn-Sham potential $V_{\rm KS}(x)$ (solid black) of the 1D cluster and the 20 occupied KS orbitals (coloured) plotted as $|\varphi_{i}(x)|^{2}+\varepsilon_{i}$.
@@
We employ the t-SURFF method Tao and Scrinzi (2012); Bauer (2017) to compute photoelectron spectra. In 1D and velocity gauge, the momentum-space amplitudes collected at the left (right) surfaces $x_{L}$ ($x_{R}$) read (for each occupied orbital $\varphi_{i}$)
@@ II.1 Numerical details
We use $N=2000$ grid points with a spacing $\Delta x=0.5$, a time step $\Delta t=0.25$, and complex absorbing boundaries at the edges. The convolution theorem is used to calculate $V_{\text{H}}[n](x,t)$. Real‑time propagation of (6) uses the Crank-Nicolson scheme with a predictor-corrector step and absorbing boundaries Bauer (2017). The initial configuration, solving (2) is obtained with imaginary-time propagation using also the Crank-Nicolson scheme. The t-SURFF surfaces are placed at $x_{L}=0.25\,N\Delta x$ and $x_{R}=0.75\,N\Delta x$. Convergence was verified with respect to box size, time step, position of the t-SURFF surfaces, and the $k$-grid used for the electron spectra.
@@ III Identification of collective modes
In order to identify whether a linear-response peak is due to single-particle transitions or collective in nature, it is useful to look at the response of individual orbitals to a perturbation. To that end we apply a weak $\delta$-kick in the electric field (i.e., a $\Theta$-step in the vector potential) to the ground state configuration. Let us first look at the Fourier-transformed total dipole
The pre-factor $\Omega^{2}$ arises because we are actually interested in the FFT of $\ddot{D}(t)$, the dipole acceleration, so that $\Omega$ is the frequency at which the dipolar kicked system radiates.
Figure 2 shows results for the “frozen” and the fully dynamic system. Here, “frozen” means that we set $V_{\text{KS}}[n](x,t)$ in (6) to $V_{\text{KS}}[n](x,0)$. In such a simulation, one expects peaks at energy differences $\varepsilon_{j}-\varepsilon_{i}$. The expected positions for such peaks are indicated by vertical dashed lines. The strongest peak is at the energy difference between the highest occupied KS orbital ($\varepsilon_{20}=-0.1709$) and the first unoccupied ($\varepsilon_{21}=-0.0984$), indicated by the first gray dashed line. There is no peak at $\varepsilon_{22}-\varepsilon_{20}$ (second gray dashed line) because of the dipole selection rule, while there is one at $\varepsilon_{23}-\varepsilon_{20}$ (third gray dashed line). The same repeats for the second-highest occupied KS orbital at $\varepsilon_{21}-\varepsilon_{19}$ (first green dashed line, no peak), $\varepsilon_{22}-\varepsilon_{19}$ (second green dashed line, peak present) etc. The yellow dashed lines indicate $\varepsilon_{21}-\varepsilon_{18}$, $\varepsilon_{22}-\varepsilon_{18}$, $\varepsilon_{23}-\varepsilon_{18}$. There are no peaks for transitions between initially populated KS levels $i,j$ because $\varepsilon_{i}\to\varepsilon_{j}$ interferes destructively with $\varepsilon_{j}\to\varepsilon_{i}$ (this is how the Pauli principle is realized in a system of non-interacting identical particles).
Figure 2: Linear-response spectra. Black: full, dynamic response according to (6). Blue: response for frozen KS potential $V_{\text{KS}}[n](x,0)$. Vertical dashed lines indicate KS level differences. Two strong peaks at $\omega_{A}=0.106$ and $\omega_{B}=0.156$ appear in the fully dynamic (“unfrozen”) response.
In the full, dynamical response (labeled “unfrozen” in Fig. 2 ) two strong peaks at
@@
are observed. In order to prove that these are collective responses we show orbital-resolved linear response spectra in Fig. 3. In such spectra, one sees all the transitions to other states, which, however, are different for the different KS orbitals. Only in the case of a harmonic-oscillator-like KS potential with equidistant levels would there be only one common response at the harmonic oscillator frequency for all orbitals. However, we do see response peaks in Fig. 3 that are at the same $\Omega$ for all KS orbitals despite the fact that $V_{\text{KS}}[n](x)$ is not a harmonic-oscillator potential. By definition, this is a collective response, as all KS orbitals oscillate in unison with this frequency. The first two of such collective responses are at $\Omega=\omega_{A}$ and $\omega_{B}$.
Figure 3: Orbital-resolved spectra. A strong response of all KS orbitals at a given $\Omega$ indicates a collective mode. Instead, orbital-dependent responses are due to single-particle transitions.
The spatial and temporal evolution of the current density for the two collective modes is shown in Fig. 4. The oscillations are plotted for times following the laser pulses that selectively excited these modes, using laser frequencies of $\omega_{L}=\omega_{A}/3$ and $\omega_{L}=\omega_{B}/3$, respectively. Although these modes could in principle also be excited by single-photon absorption, we focus in this work on photoelectron emission following nonlinear plasmon excitation. Since $2\omega_{L}$ excitations are forbidden by selection rules, the lowest-order nonlinear plasmon excitations occur at $3\omega_{L}$. As expected, Fig. 4 shows that the lower-energy mode $\omega_{A}$ corresponds to a collective oscillation of the entire electron density, while the higher-energy mode exhibits two nodes near the edges of the one-dimensional cluster. Consequently, the current density inside the cluster oscillates in the opposite direction to the surface currents.
Figure 4: Oscillating current densities for the collective modes $\omega_{A}=0.106$ (upper panel) and $\omega_{B}=0.156$ (lower panel) as a function of time (in units of $2\pi/\omega_{L}$) and space. The modes were excited with laser pulses of frequencies $\omega_{L}=\omega_{A}/3$ and $\omega_{L}=\omega_{B}/3$, respectively. The oscillating current densities are shown for times after the laser pulses.
@@ IV Photoelectron spectra
Figure 5 shows the photoelectron spectrum (12) for a laser pulse with parameters $a_{0}=0.004$, $\omega_{L}=0.052=\omega_{B}/3$, and $N_{\text{cyc}}=20$. Also shown are the individual spectra for the highest occupied KS orbitals $i=20,19,18$. Lower-lying KS orbitals do not play a role for the total electron yield. Besides the expected ATI peaks, separated by $\omega_{L}$, there are very sharp spiky features. The more detailed analysis presented below shows that these features arise from plasmon excitations that decay too slowly in the TDDFT treatment. Extended post-pulse propagation within the t-SURFF framework is required to collect low-energy electrons at the distant t-SURFF surfaces. During this propagation, t-SURFF also accumulates probability amplitude associated with plasmon-assisted electron emission occurring after the laser pulse. The longer the post-pulse duration, the more pronounced and sharper these peaks become.
@@
Figure 5: Total photoelectron spectrum (red) and the individual contributions from the three highest initially populated KS orbitals for a pulse that dominantly excites $\omega_{B}=0.156$ with three laser photons. Narrow spikes are equally spaced by $\omega_{B}$. Laser parameters are $a_{0}=0.004$, $\omega_{L}=0.052=\omega_{B}/3$, $N_{\text{cyc}}=20$.
@@ IV.1 Time-of-flight analysis
Figure 6: Windowed t-SURFF spectrum as a function of time (in units of $2\pi/\omega_{L}$) and energy $E=k^{2}/2$. Laser parameters are the same as in Fig. 5.
Figure 6 shows the windowed t-SURFF spectrum for the same laser parameters as in Fig. 5. The width of the Gaussian time-window was one laser cycle so that the ATI peaks due to direct, laser-induced ionization are clearly visible. All ATI-emitted electrons arrive before $\omega_{L}t/2\pi\simeq 30$ at the t-SURFF surfaces, which is consistent with them being emitted during the laser pulse plus their time-of-flight. After the ATI peaks, the plasmon-assisted emission builds up and is clearly visible as horizontal stripes at positions $E^{(A,B)}_{in}=\varepsilon_{i}+n\,\omega_{A,B}$. The stripes are broader in energy than the narrow spikes in the total spectrum in Fig. 5 because of the narrow time-window. This also shows that the plasmon-assisted electron emission peaks would be broader (and probably closer to reality) if we stop the simulation soon after the laser pulse. However, at that time the low-energetic electrons have not yet arrived at the t-SURFF surfaces and thus are missing in the spectra. We also experimented with “freezing” the KS potential after the laser pulse, which, however, yields other artefacts.